A carpenter works one third as long after lunch as she worked before lunch. If she works a total of 8 hrs, how long did she works before lunchs
Found 3 solutions by ewatrrr, stanbon, advanced_Learner:
Answer by ewatrrr(24785)

x hours worked before lunch
x/3 + x = 8
(4/3)x = 8
x = 8(3/4)
x = 6hr
And
2hr + 6hr = 8hr CHECKS!

Answer by stanbon(75887)

A carpenter works one third as long after lunch as she worked before lunch. If she works a total of 8 hrs, how long did she work before lunch?
--------------------
Before lunch:: 3x
After lunch:: x
-----
Equation:
3x + x = 8 hrs
4x = 8
x = 2
----
Before lunch:: 3x = 3*2 = 6 hrs
After lunch:: x = 2 hrs
------------------
Cheers,
Stan H.
-------------------

Answer by advanced_Learner(501)

A carpenter works one third as long after lunch as she worked before lunch. If she works a total of 8 hrs, how long did she works before lunchs

before lunch x
after 1/3x
x+1/3x=8
3x+x=24
4x=24
x=6
before lunch 6 hours
after lunch 2 hours
total 8 hours.